# An interpretable family of projected normal distributions and a related copula model for Bayesian analysis of hypertoroidal data

### v1

@@ 2.2 Definition and basic properties
Figure 1: Plot of the density (9) for: (a) $\kappa_{1}=0.49$ with $\mu_{1}=0$ (solid), $\mu_{1}=\pi/3$ (dashed), $\mu_{1}=2\pi/3$ (dotted), and $\mu_{1}=-2\pi/3$ (dot-dashed); and (b) $\mu_{1}=0$ with $\kappa_{1}=0$ (dot-dashed), $\kappa_{1}=0.49$ (dotted), $\kappa_{1}=1.1$ (dashed), and $\kappa_{1}=2.45$ (solid).
@@
       Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Contour plots of density (12) of $(\Theta_{1},\Theta_{2})$ for: (a) $\mu_{1}=\mu_{2}=0$, $\kappa_{2}=0.49$, with various combinations of $(\kappa_{1},\rho_{12})$; and (b) the same parameters as in the upper next frame of (a), but with the following adjustments: (i) $\mu_{1}=0$ is replaced by $\mu_{1}=\pi/2$, (ii) $\mu_{2}=0$ is replaced by $\mu_{2}=-\pi/2$, (iii) $\rho_{12}=0.9$ is replaced by $\rho_{12}=-0.9$, and (iv) $\kappa_{2}=0.49$ is replaced by $\kappa_{2}=2.45$.
Figure 1 shows the univariate marginal density (9) of $\Theta_{1}$ for selected values of the parameter. The four values of $\kappa_{1}$ used in Figure 1, namely, $(0,0.49,1.1,2.45)$, approximately correspond to mean resultant lengths $|E(e^{{i}\Theta_{1}})|$ of $(0,0.3,0.6,0.9)$, respectively. The mean resultant length is a measure of concentration for circular random variable, ranging from zero for uniformity to one for a point mass (see, e.g., Mardia and Jupp,, 1999, Section 3.4.2). As seen in Figure 1(a), the parameter $\mu_{1}$ represents the mode of the univariate density (9). The density (9) reaches its maximum value at $\theta_{1}=\mu_{1}$ and its minimum value at $\theta_{1}=\mu_{1}+\pi$. It is also clear that the density (9) is symmetric about $\theta_{1}=\mu_{1}$. Figure 1(b) shows how parameter $\kappa_{1}$ controls the concentration of the univariate marginal distribution. If $\kappa_{1}=0$, then the marginal distribution of $\Theta_{1}$ is uniform on the circle. The greater the value of $\kappa_{1}$, the greater the concentration of the marginal density of $\Theta_{1}$ around $\theta_{1}=\mu_{1}$. It can be shown that, as $\kappa_{1}\to\infty$, the marginal distribution of $\Theta_{1}$ converges to a point distribution with a singularity at $\theta_{1}=\mu_{1}$.
@@
Figure 2 displays the bivariate marginal density of TPN of $(\Theta_{1},\Theta_{2})$, whose functional form will be given later in (12), for selected combinations of the parameters. In Figure 2(a), for fixed $\kappa_{1}$, it can be seen that the parameter $\rho_{12}$ controls the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$. When $\rho_{12}=0$, $\Theta_{1}$ and $\Theta_{2}$ are independent. As $|\rho_{12}|$ increases, the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$ also increases. The comparison between frame (b)(iii) and the frame directly above it in (a) shows that positive and negative values of $\rho_{12}$ imply positive and negative associations between the two variables, respectively. The role of the parameter $\kappa_{1}$ is evident from Figure 2(a) for fixed $\rho_{12}$, enhancing the interpretation of $\kappa_{1}$ for the univariate marginal given in Figure 1(b). When $\kappa_{1}=0$, the TPN is highly dispersed with respect to $\theta_{1}$. As $\kappa_{1}$ increases, the density becomes more concentrated around $\theta_{1}=\mu_{1}$. The parameter $\kappa_{2}$ plays an analogous role with respect to $\theta_{2}$; see frame (iv) of Figure 2(b). Finally, frames (i) and (ii) of Figure 2(b), in comparison with the frames directly above them, suggest that the joint density achieves its maximum at $(\theta_{1},\theta_{2})=(\mu_{1},\mu_{2})$.
2.3 Probability density function
Although the probability density function of the TPN cannot be written in closed form, here we show some new results that are also useful for the definition of the copula model based of the TPN. Let $\bm{R}=(R_{1},\ldots,R_{d})^{\top}$, then it is straightforward to see that the joint density of $(\bm{R},\bm{\Theta})$ can be expressed as
@@ 2.3 Probability density function
where $\bm{x}_{c}$ and $\bm{x}_{s}$ in (10) must be seen as functions of $(\bm{r},\bm{\theta})$, as defined in equation (8). It follows that the density of $\bm{\Theta}$ is
@@
As is the case for the PN distribution on the hypersphere, it is difficult to obtain a closed form expression for $f(\bm{\theta})$ in the general case, but we will show that under our parametrization, it is possible to do it in some special case, see Section 3, where the copula model based on the TPN is introduced. Using a method similar to the one in Mastrantonio, (2018), this form of the density will be useful for the implementation of the MCMC algorithm for both the TPN and the copula case, see Section 5.
The density (11) has an alternative expression which is given below.
@@
$\begin{split}f(\bm{\theta})=&\ \frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\exp\left[-\frac{1}{2}\bm{\kappa}^{\top}\left(\mathbf{I}-\bm{\Sigma}_{c}^{-1}\mathbf{T}_{c}\mathbf{V}\mathbf{T}_{c}\right)\bm{\Sigma}_{c}^{-1}\bm{\kappa}\right]\\ &\times\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\mathbf{V}^{-1}(\bm{r}-\bm{\eta})\right]\prod_{j=1}^{d}r_{j}\,d\bm{r},\end{split}$ (12)
where $\bm{\eta}=\mathbf{V}\mathbf{T}_{c}\bm{\Sigma}_{c}^{-1}\bm{\kappa}$, $\mathbf{V}=(\mathbf{T}_{c}\bm{\Sigma}_{c}^{-1}\mathbf{T}_{c}+\mathbf{T}_{s}\bm{\Sigma}^{-1}\mathbf{T}_{s})^{-1}$, $\mathbf{T}_{c}=\mbox{\rm diag}(\cos(\theta_{1}-\mu_{1}),\ldots,\cos(\theta_{d}-\mu_{d}))$, and $\mathbf{T}_{s}=\mbox{\rm diag}(\sin(\theta_{1}-\mu_{1}),\ldots,\sin(\theta_{d}-\mu_{d}))$.
Since the exponential inside the integral of (12) corresponds to the kernel of a multivariate truncated normal distribution with parameters $\{\bm{\eta},\mathbf{V}\}$ for the variable $\bm{r}$, the integral is equal to
$E(R_{1}R_{2}\dots R_{d})\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\mathbf{V}^{-1}(\bm{r}-\bm{\eta})\right]d\bm{r},$ (13)
where $E(R_{1}R_{2}\dots R_{d})$ is the expectation with respect to the multivariate truncated normal distribution on the positive orthant, and the integral in (13) is the normalizing constant of the truncated normal density. Although $E(R_{1}R_{2}\dots R_{d})$ cannot be computed in closed form, numerical methods exist that allow to obtain its value, see Tallis, (1961), Lee, (1983) and Leppard and Tallis, (2018). The integral part is proportional to the value of the cumulative distribution function of the multivariate normal distribution $N_{d}(-\bm{\eta},\mathbf{V})$ at the origin, and can be computed straightforwardly using numerical methods.
As we can see from the following theorem, there are condition in which the matrix $\mathbf{V}$ in Theorem 4 can be greatly simplified:
@@
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\mathbf{D}\bm{\Sigma}\mathbf{D}$, where $\mathbf{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Then $\mathbf{V}$ is of the form
$\mathbf{V}=\begin{pmatrix}s_{11}&s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&\cdots&s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})\\ s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&s_{22}&\cdots&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})\\ \vdots&\vdots&\ddots&\vdots\\ s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})&\cdots&s_{dd}\end{pmatrix}^{-1},$
@@
Then there always exists $(\delta_{j},\delta_{k})$ such that $\bm{\Sigma}^{(j,k)}_{c}=\mathbf{D}\bm{\Sigma}^{(j,k)}\mathbf{D}$. In fact, by assuming $\delta_{j}=1$ and $\delta_{k}=\text{sign}(\rho_{jk})$, it is easy to see that $\bm{\Sigma}^{(j,k)}_{c}=\mathbf{D}\bm{\Sigma}^{(j,k)}\mathbf{D}$ holds for any $\rho_{jk}$, where $\mbox{sign}(\rho)=1$ if $\rho\geq 0$ and $\mbox{sign}(\rho)=-1$ if $\rho<0$. Thus Theorems 2 and 4 imply that the bivariate marginal density of $(\Theta_{j},\Theta_{k})$ can be expressed in the same form as (12) with $d=2$, and Corollary 1 suggests that the matrix $\mathbf{V}$ which corresponds to the bivariate case is given by
$\mathbf{V}$ $=\left\{\frac{1}{1-\rho_{jk}^{2}}\begin{pmatrix}1&\tau(\theta_{j},\theta_{k})\\ \tau(\theta_{j},\theta_{k})&1\end{pmatrix}\right\}^{-1}=\frac{1-\rho_{jk}^{2}}{1-\tau(\theta_{j},\theta_{k})^{2}}\begin{pmatrix}1&-\tau(\theta_{j},\theta_{k})\\ -\tau(\theta_{j},\theta_{k})&1\end{pmatrix}$, (14)
@@ 3.1 The copula case of the TPN distribution
Let a random vector $\bm{\Theta}=(\Theta_{1},\ldots,\Theta_{d})^{\top}$ follow $TPN_{d}(\bm{\mu},\bm{0},\bm{\Sigma})$. Then, from (11) and (12), it follows that the density of $\bm{\Theta}$ reduces to
$\begin{split}c(\bm{\theta})&=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}\left(\bm{x}_{c}^{\top}\bm{\Sigma}_{c}^{-1}\bm{x}_{c}+\bm{x}_{s}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s}\right)\right]\prod_{j=1}^{d}r_{j}d\bm{r}\\ &=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left(-\frac{\bm{r}^{\top}\mathbf{V}^{-1}\bm{r}}{2}\right)\prod_{j=1}^{d}r_{j}\,d\bm{r}.\end{split}$ (16)
It should be noted that, in the second expression of the density (16), the variable $\bm{\theta}$ appears only inside the matrix $\mathbf{V}$. Then, under the assumption in Corollary 1, the density (16) is a function of $(\delta_{i}\bar{\theta}_{i}-\delta_{j}\bar{\theta}_{j})_{1\leq i,j\leq d}$.
It immediately follows from Theorem 3 that each univariate marginal of this submodel is the uniform distribution on the circle with density
@@
This means that (16) is the density of a hypertoroidal copula model. Following standard practice in copula theory, we denote the density of the copula case of our family by $c$.
@@
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\mathbf{D}\bm{\Sigma}\mathbf{D}$, where $\mathbf{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Let $s_{jk}<0$ for $1\leq j,k\leq d,\ j\neq k$, where $s_{jk}$ denotes the $(j,k)$ element of $\bm{\Sigma}_{c}^{-1}$. Then the copula density (16) takes its maximum values if and only if $\delta_{1}(\theta_{1}-\mu_{1})=\delta_{2}(\theta_{2}-\mu_{2})=\cdots=\delta_{d}(\theta_{d}-\mu_{k})$.
Note that the assumption of this theorem always holds when $d=2$. Although this is not the case for $d\geq 3$ in general, our simulation results suggest that the condition of the theorem often holds for $d=3$ for randomly generated $\bm{\Sigma}$ matrices with positive off-diagonal elements.
@@
$\begin{split}c_{2}(\theta_{j},\theta_{k})=&\ \frac{1}{(2\pi)^{2}}\frac{1-\rho^{2}}{\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{3/2}}\\ &\times\left\{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}-\tau(\theta_{j},\theta_{k})\,\text{atan}^{*}\left(\frac{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}}{\tau(\theta_{j},\theta_{k})}\right)\right\},\end{split}$ (18)
where $\tau(\theta_{i},\theta_{j})$ is given in (15) and $\text{atan}^{*}$ is defined in Section 2.1.
This theorem implies that the distribution (16) can be considered a multivariate extension of the bivariate copula model of Wehrly and Johnson, (1980), which is defined by a density that depends only on $\theta_{1}-\theta_{2}$ or $\theta_{1}+\theta_{2}$. The model of Wehrly and Johnson, (1980) has played an important role in directional statistics as a general model for bivariate circular data; see Jones et al., (2015) for detailed properties of their model. Note that the closed-form expression (18) of the bivariate marginal comes from the fact that the conditions of Corollary 1 are always met in this case. It is also possible to avoid the use of $\text{atan}^{*}$ and adopt the ordinary arctangent function which takes values on $\pi/2,\pi/2]$. This can be done by using the relationship $\text{atan}^{*}[\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}/\tau(\theta_{j},\theta_{k})]=\pi/2-\arctan[\tau(\theta_{j},\theta_{k})/\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}]$, which gives us an alternative expression for the density (18):
@@
$\times\left[\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}+\tau(\theta_{j},\theta_{k})\left\{\arctan\left(\frac{\tau(\theta_{j},\theta_{k})}{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}}\right)-\frac{\pi}{2}\right\}\right].$

### v2

@@ 2.2 Definition and basic properties
Figure 1: Plots of the univariate marginal density (5) of $\Theta_{1}$ for: (a) $\kappa_{1}=0.49$ with four values of $\mu_{1}$; and (b) $\mu_{1}=0$ with four values of $\kappa_{1}$.
Figure 1 shows the univariate marginal density (5) of $\Theta_{1}$ for selected values of the parameter. The four values of $\kappa_{1}$ used in Fig. 1, namely, $(0,0.49,1.1,2.45)$, approximately correspond to mean resultant lengths $|E(e^{{i}\Theta_{1}})|$ of $(0,0.3,0.6,0.9)$, respectively. The mean resultant length is a measure of concentration for circular random variable, ranging from zero for uniformity to one for a point mass (see, e.g., Mardia1999a, §3.4.2). As seen in Fig. 1(a), the parameter $\mu_{1}$ represents the mode of the univariate density (5). The density (5) reaches its maximum value at $\theta_{1}=\mu_{1}$ and its minimum value at $\theta_{1}=\mu_{1}+\pi$. Also the density (5) is symmetric about $\theta_{1}=\mu_{1}$. Figure 1(b) shows how parameter $\kappa_{1}$ controls the concentration of the univariate marginal distribution. If $\kappa_{1}=0$, then the marginal distribution of $\Theta_{1}$ is uniform on the circle. The greater the value of $\kappa_{1}$, the greater the concentration of the marginal density of $\Theta_{1}$ around $\theta_{1}=\mu_{1}$. As $\kappa_{1}\to\infty$, the marginal distribution of $\Theta_{1}$ converges to a point distribution with a singularity at $\theta_{1}=\mu_{1}$.
@@
Figure 2: Contour plots of the marginal density (10) of $(\Theta_{1},\Theta_{2})$ for: (a) $\mu_{1}=\mu_{2}=0$, $\kappa_{2}=0.49$, with various combinations of $(\kappa_{1},\rho_{12})$; and (b) the same parameters as in the upper next frame of (a), but with the following adjustments: (i) $\mu_{1}=0$ is replaced by $\mu_{1}=\pi/2$, (ii) $\mu_{2}=0$ is replaced by $\mu_{2}=-\pi/2$, (iii) $\rho_{12}=0.9$ is replaced by $\rho_{12}=-0.9$, and (iv) $\kappa_{2}=0.49$ is replaced by $\kappa_{2}=2.45$.
@@
Figure 2 displays the bivariate marginal density of $(\Theta_{1},\Theta_{2})$, whose functional form will be given later in (10), for selected combinations of the parameters. In Fig. 2(a), for fixed $\kappa_{1}$, it can be seen that the parameter $\rho_{12}$ controls the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$. When $\rho_{12}=0$, $\Theta_{1}$ and $\Theta_{2}$ are independent. As $|\rho_{12}|$ increases, the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$ also increases. The comparison between frame (b)(iii) and the frame directly above it in (a) shows that positive and negative values of $\rho_{12}$ imply positive and negative associations between the two variables, respectively. The role of the parameter $\kappa_{1}$ is evident from Fig. 2(a) for fixed $\rho_{12}$, enhancing the interpretation of $\kappa_{1}$ for the univariate marginal given in Fig. 1(b). When $\kappa_{1}=0$, the TPN is highly dispersed with respect to $\theta_{1}$. As $\kappa_{1}$ increases, the density becomes more concentrated around $\theta_{1}=\mu_{1}$. The parameter $\kappa_{2}$ plays an analogous role with respect to $\theta_{2}$; see frame (iv) of Fig. 2(b). Finally, frames (i) and (ii) of Fig. 2(b), in comparison with the frames directly above them, suggest that the joint density achieves its maximum at $(\theta_{1},\theta_{2})=(\mu_{1},\mu_{2})$.
2.3 Form of probability density function
Introducing latent variables $R_{1},\ldots,R_{d}>0$, the following relationship holds between $\bm{\Theta}$ and $(\bm{X}_{c},\bm{X}_{s})$ in Definition 1:
$\bm{X}_{c}=\begin{pmatrix}R_{1}\cos(\Theta_{1}-\mu_{1})\\ R_{2}\cos(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\cos(\Theta_{d}-\mu_{d})\end{pmatrix}\quad\mbox{and}\quad\bm{X}_{s}=\begin{pmatrix}R_{1}\sin(\Theta_{1}-\mu_{1})\\ R_{2}\sin(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\sin(\Theta_{d}-\mu_{d})\end{pmatrix}$. (6)
Let $\bm{R}=(R_{1},\ldots,R_{d})^{\top}$. Then the joint density of $(\bm{R},\bm{\Theta})$ can be expressed as
@@ 2.3 Form of probability density function
$\hskip 227.62204pt(\bm{r}\in(0,\infty)^{d},\ \bm{\theta}\in[-\pi,\pi)^{d})$ (8)
where $\bm{x}_{c}$ and $\bm{x}_{s}$ in (7) should be regarded as functions of $(\bm{r},\bm{\theta})$, as defined in (6). It follows that the density of $\bm{\Theta}$ is
@@
As with existing distributions based on projections (Wang2013a; Wang2014; Hernandez-Stumpfhauser2017; Mastrantonio2018), it is also difficult to obtain a closed-form expression for (7) in the general case, apart from a special case discussed in §3. However, by treating $\bm{r}$ as a latent variable, the form of our density (7) becomes useful for simplifying the MCMC algorithm; see §5.
@@
$\begin{split}f(\bm{\theta})=&\ \frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\exp\left[-\frac{1}{2}\bm{\kappa}^{\top}\left(\bm{I}-\bm{\Sigma}_{c}^{-1}\bm{T}_{c}\bm{V}\bm{T}_{c}\right)\bm{\Sigma}_{c}^{-1}\bm{\kappa}\right]\\ &\times\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\bm{V}^{-1}(\bm{r}-\bm{\eta})\right]\prod_{j=1}^{d}r_{j}\,d\bm{r},\end{split}$ (10)
where $\bm{\eta}=\bm{V}\bm{T}_{c}\bm{\Sigma}_{c}^{-1}\bm{\kappa}$, $\bm{V}=(\bm{T}_{c}\bm{\Sigma}_{c}^{-1}\bm{T}_{c}+\bm{T}_{s}\bm{\Sigma}^{-1}\bm{T}_{s})^{-1}$, $\bm{T}_{c}=\mbox{\rm diag}\{\cos(\theta_{1}-\mu_{1}),\ldots,\cos(\theta_{d}-\mu_{d})\}$, and $\bm{T}_{s}=\mbox{\rm diag}\{\sin(\theta_{1}-\mu_{1}),\ldots,\sin(\theta_{d}-\mu_{d})\}$.
Since the exponential inside the integral of (10) corresponds to the kernel of a multivariate truncated normal distribution with parameters $\{\bm{\eta},\bm{V}\}$ for the variable $\bm{r}$, the integral is equal to
$E(R_{1}R_{2}\cdots R_{d})\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\bm{V}^{-1}(\bm{r}-\bm{\eta})\right]d\bm{r},$ (11)
where $E(R_{1}R_{2}\cdots R_{d})$ is the expectation with respect to the multivariate truncated normal distribution on the positive orthant, and the integral in (11) is the normalizing constant of the truncated normal density. Although $E(R_{1}R_{2}\cdots R_{d})$ cannot be computed in closed form, numerical methods exist that allow to obtain its value; see Tallis, LEE1983245 and Leppard. The integral part is proportional to the value of the cumulative distribution function of the multivariate normal distribution ${\cal N}_{d}(-\bm{\eta},\bm{V})$ at the origin, and can be computed straightforwardly using numerical methods.
@@
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\bm{D}\bm{\Sigma}\bm{D}$, where $\bm{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Then $\bm{V}$ is of the form
$\bm{V}=\begin{pmatrix}s_{11}&s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&\cdots&s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})\\ s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&s_{22}&\cdots&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})\\ \vdots&\vdots&\ddots&\vdots\\ s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})&\cdots&s_{dd}\end{pmatrix}^{-1},$
@@
Then there always exists $(\delta_{j},\delta_{k})$ such that $\bm{\Sigma}^{(j,k)}_{c}=\bm{D}\bm{\Sigma}^{(j,k)}\bm{D}$. In fact, by assuming $\delta_{j}=1$ and $\delta_{k}=\text{sign}(\rho_{jk})$, it is easy to see that $\bm{\Sigma}^{(j,k)}_{c}=\bm{D}\bm{\Sigma}^{(j,k)}\bm{D}$ holds for any $\rho_{jk}$, where $\mbox{sign}(\rho)=1$ if $\rho\geq 0$ and $\mbox{sign}(\rho)=-1$ if $\rho<0$. Thus Theorems 2 and 4 imply that the bivariate marginal density of $(\Theta_{j},\Theta_{k})$ can be expressed in the same form as (10) with $d=2$, and Corollary 1 suggests that the matrix $\bm{V}$ which corresponds to the bivariate case is
$\bm{V}$ $=\left\{\frac{1}{1-\rho_{jk}^{2}}\begin{pmatrix}1&\tau(\theta_{j},\theta_{k})\\ \tau(\theta_{j},\theta_{k})&1\end{pmatrix}\right\}^{-1}=\frac{1-\rho_{jk}^{2}}{1-\tau(\theta_{j},\theta_{k})^{2}}\begin{pmatrix}1&-\tau(\theta_{j},\theta_{k})\\ -\tau(\theta_{j},\theta_{k})&1\end{pmatrix}$, (12)
@@ 3.1 The copula case of the TPN distribution
Let a random vector $\bm{\Theta}=(\Theta_{1},\ldots,\Theta_{d})^{\top}$ follow $\mathrm{TPN}_{d}(\bm{\mu},\bm{0},\bm{\Sigma})$. Then it follows from (9) and (10) that the density of $\bm{\Theta}$ reduces to
$\begin{split}c(\bm{\theta})&=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}\left(\bm{x}_{c}^{\top}\bm{\Sigma}_{c}^{-1}\bm{x}_{c}+\bm{x}_{s}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s}\right)\right]\prod_{j=1}^{d}r_{j}d\bm{r}\\ &=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left(-\frac{\bm{r}^{\top}\bm{V}^{-1}\bm{r}}{2}\right)\prod_{j=1}^{d}r_{j}\,d\bm{r},\end{split}$ (14)
where $\bm{V}$ is as in Theorem 4. In the second expression of the density (14), the variable $\bm{\theta}$ appears only inside the matrix $\bm{V}$. Then, under the assumption in Corollary 1, the density (14) is a function of $(\delta_{i}\bar{\theta}_{i}-\delta_{j}\bar{\theta}_{j})_{1\leq i,j\leq d}$.
Theorem 3 implies that each univariate marginal of this submodel is the uniform distribution on the circle with density
@@
This means that (14) is the density of a hypertoroidal copula model. We denote by $c$ the density in the copula case of our family, and by $c_{j}$ its $j$-variate marginal density.
@@
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\bm{D}\bm{\Sigma}\bm{D}$, where $\bm{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Let $s_{jk}<0$ for $1\leq j,k\leq d,\ j\neq k$, where $s_{jk}$ is as in Corollary 1. Then the copula density (14) takes its maximum values if and only if $\delta_{1}(\theta_{1}-\mu_{1})=\delta_{2}(\theta_{2}-\mu_{2})=\cdots=\delta_{d}(\theta_{d}-\mu_{k})$.
The assumption of this theorem always holds when $d=2$. Although this is not the case for $d\geq 3$ in general, our simulation results suggest that the condition of the theorem often holds for $d=3$ for randomly generated $\bm{\Sigma}$ matrices with positive off-diagonal elements.
@@
$\begin{split}c_{2}(\theta_{j},\theta_{k})=&\ \frac{1}{(2\pi)^{2}}\frac{1-\rho^{2}}{\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{3/2}}\\ &\times\left(\left\{1-\tau(\theta_{j},\theta_{k})^{2}\right\}^{1/2}-\tau(\theta_{j},\theta_{k})\,\text{atan}^{*}\left[\frac{\left\{1-\tau(\theta_{j},\theta_{k})^{2}\right\}^{1/2}}{\tau(\theta_{j},\theta_{k})}\right]\right),\end{split}$ (16)
where $\tau(\theta_{i},\theta_{j})$ and $\text{atan}^{*}$ are as in (13) and §2.1, respectively.
This theorem implies that the distribution (14) can be considered a multivariate extension of the bivariate copula model of Wehrly1980 whose density depends only on $\theta_{1}-\theta_{2}$ or $\theta_{1}+\theta_{2}$. The model of Wehrly1980 has played an important role in directional statistics as a general model for bivariate circular data; see Jones2015 for its detailed properties. The closed-form expression (16) of the bivariate marginal comes from the fact that the conditions of Corollary 1 are always met in this case. It is also possible to avoid the use of $\text{atan}^{*}$ and adopt the ordinary arctangent function which takes values on $\pi/2,\pi/2]$. This can be done by using the relationship $\text{atan}^{*}[\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}/\tau(\theta_{j},\theta_{k})]=\pi/2-\arctan[\tau(\theta_{j},\theta_{k})/\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}]$, which gives us an alternative expression for the density (16):
@@
$\times\left(\left\{1-\tau(\theta_{j},\theta_{k})^{2}\right\}^{1/2}+\tau(\theta_{j},\theta_{k})\left[\arctan\left(\frac{\tau(\theta_{j},\theta_{k})}{\left\{1-\tau(\theta_{j},\theta_{k})^{2}\right\}^{1/2}}\right)-\frac{\pi}{2}\right]\right).$